For the function f(x)=(1/4)x+2, find the value of x for which f(x)=x
Answer by nerdybill(7384)

For the function f(x)=(1/4)x+2, find the value of x for which f(x)=x
.
set f(x) to x and solve for x:
f(x)=(1/4)x+2
x=(1/4)x+2
x-(1/4)x = 2
x(1- 1/4) = 2
x(4/4 - 1/4) = 2
x(3/4) = 2
x = 2(4/3)
x = 8/3